The following appears to prove that any two numbers are equal. Obviously it is wrong; can you spot the flaw?
Let a and b be any two different numbers. Define x as the difference between them:
x = b - a
Multiply both sides of the equation by (b-a):
x(b – a) = (b – a)(b – a)
bx – ax = b^2 – 2ab + a^2
Add –bx+ab-a^2 to both sides and simplify:
bx – ax – bx + ab – a^2 = b^2 – 2ab + a^2 – bx + ab – a^2
bx – bx – ax + ab – a^2 = –bx + b^2 – 2ab + ab + a^2 – a^2
–ax + ab – a^2 = –bx + b^2 – ab
Both sides of this equation have a common factor:
a(–x + b – a) = b(–x + b – a)
Divide both sides by (-x+b-a):
a(–x + b – a)/-x + b-a = b(–x + b – a)/ -x + b - a
a = b

What was my mistake?

Answer by Edwin McCravy(20054)


The following appears to prove that any two numbers are equal.
Obviously it is wrong; can you spot the flaw? 

Let a and b be any two different numbers. Define x as the 
difference between them:

x = b - a
Multiply both sides of the equation by (b-a):
x(b – a) = (b – a)(b – a)
bx – ax = b^2 – 2ab + a^2
Add –bx+ab-a^2 to both sides and simplify:
bx – ax – bx + ab – a^2 = b^2 – 2ab + a^2 – bx + ab – a^2
bx – bx – ax + ab – a^2 = –bx + b^2 – 2ab + ab + a^2 – a^2
–ax + ab – a^2 = –bx + b^2 – ab
Both sides of this equation have a common factor:
a(–x + b – a) = b(–x + b – a)
Divide both sides by (-x+b-a):
a(–x + b – a)/-x + b-a = b(–x + b – a)/ -x + b - a
a = b   
 
What was my mistake?

---------------------------------------------

The step in red is the mistake.  
You're dividing by zero there. That's because -x+b-a = 0.

How do we know that -x+b-a = 0?

From the original assumption x = b - a which clearly
shows that -x+b-a = 0

Take the original assumption:

x = b - a 

Multiply both sides by -1

-x = -b + a

Add b - a to both sides:

-x+b-a = 0

Therefore -x+b-a equals 0 by the original 
assumption and you can't divide by 0.

Edwin